R. Bachimanchi, G. Ciovati, A. Freyberger, C. Hovater, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Powers, T. Reilly, C. Reece, B. Rimmer
Agenda and Notes
1. Operations Status and Summary
C50-12 installed. Testing tonight Commissioning and checkout ???.
microphonics??? Tuner dampers tested Thursday: cavities 1 6 7 8 have been damped others still show fluctuations. Tested again tonight.
Field Emission and Radiation Measurements. accumulation downstream. Cavity-cavity and module to module.
Q vs. E on the NL12. Good idea to continue!
C100 He processing. NL23 improvments on field emitters. 6 FE cavities processed. Continuing.
3. C75 Mock-up Results (Gigi/Bob)
First look at data
4. Follow up Field emission Workshop (Geng)
Indigo Site available The indico web site for the followup meeting is
5. LLRF Firmware Recs Summary (Rama)
6. Regular Reporting